# Improving and Evaluating Open Deep Research Agents

### v1

@@ 5  Open Deep Research Plus (ODR+)
This module focuses on addressing each of the sub-question identified in module (1) and is shown in lines 10–30 of Algorithm 1. The iterative process continues until all sub-questions are addressed, or some other stopping criteria is met (e.g., permissible run-time, denoted $T_{max}$, is exceeded; or a maximum number of sub-questions, denoted $D_{max}$ is exceeded).
@@
At the beginning of each iteration, the system selects an unresolved sub-question from S.subquestions (Line 12) and uses it directly as a web search query. We observed that web browsers returned a different page ranking each time the same query was submitted and therefore, to reduce variability, we submitted the same query $N_{\text{query}}$ times using webSearch (Line 13). From each of the $N_{\text{query}}$ search attempts, the top-ranked URLs are collected. The frequency of each URL across all attempts is then tallied, and the $k$ most frequently occurring URLs are selected using selectMostFrequent (Line 14) for further processing.
@@
Next, an extraction prompt extractionPrompt is constructed by invoking createExtractionPrompt with the prompt template $P_{3}$, the original user query, and the previously extracted constraints (Line 15). It is designed to instruct the LLM to extract only the portions of the page content that are relevant to the sub-question and consistent with the specified constraints. The prompt is passed, along with the full text of each selected URL, to the LLM. The model is invoked once per URL and typically returns one or more short spans of relevant text. These outputs are stored as structured findings—each consisting of the extracted text and its source URL—and are appended to S.findings, the list of accumulated findings maintained in the internal research state (Line 17).
@@
After collecting new findings, ODR+ invokes the language model using analyzeEvidence with prompt $P_{4}$, the current sub-question, and the full set of accumulated findings (Line 18). The prompt instructs the model to produce a structured response in JSON format, including fields such as a proposed answer to the sub-question (if one can be determined), a confidence score, a list of satisfied constraints, and any suggested follow-up sub-questions. The JSON output is parsed, and if a valid answer is provided, it is added to S.subAnswers (Lines 19–20). If follow-up sub-questions are suggested, they are appended to S.subquestions (Lines 21–22). Additionally, the control flow logic uses the model’s analysis—specifically the confidence score and recommendation on whether to continue—to decide whether to proceed to the next iteration or terminate the loop early (Lines 23–24).
@@
By iteratively resolving sub-questions in a goal-directed manner and updating its internal state at each step, ODR+ constructs a coherent body of evidence grounded in retrieved content, enabling robust and explainable research over complex factual queries.
Response Synthesis
The third and final module of the ODR+ system—implemented in lines 32–33 of Algorithm 1—is responsible for synthesizing the final structured answer once the iterative search loop terminates.
On line 32, the system invokes the language model using the engineered prompt $P_{5}$, along with the original user question (userQuery), the extracted constraints (constraints), and the accumulated evidence (S.findings). The prompt instructs the model to generate a final answer grounded strictly in this structured content, explicitly forbidding reliance on prior knowledge. The model is asked to produce a response in the standardized BrowseComp format:
@@
The prompt also instructs the model to compute a confidence score based on the number of key constraints satisfied by the proposed answer, relative to the total number of extracted constraints.
On line 33, the model’s output is stored in structuredResponse. The system then validates this response to ensure that all required fields are present and correctly formatted. If any field (such as the explanation, exact answer, or confidence score) is missing or malformed, fallback values are inserted. For example, the system may assign "Unknown" as the answer and a default confidence score of 10%.
This validation step ensures that every final output is complete, properly structured, and ready for automated evaluation. By grounding all answers in retrieved evidence and enforcing a consistent output format, ODR+ produces responses that are reliable, transparent, and evaluation-ready.
@@ 6  Numerical Experiments
We conduct experiments on our BrowseComp-Small benchmark (see Sec. 3), which comprises two disjoint sets of sixty questions: a training set, and a testing set. We evaluate several competing DRA systems on the sixty test questions: ODR, ODR+, Claude-DeepResearch (Anthropic) [15], and Gemini-DeepResearch (Google 2025) [16]. We used the sixty questions in the BrowseComp-Small training set to develop ODR+.
ODR+ Development and Hyperparameter Settings
All development of ODR+ was done using the sixty training questions in our BrowseComp-Small benchmark. This was done to minimize the potential of overfitting the design of ODR+ to the testing questions. Many steps of ODR+ (and ODR) utilize an LLM, and we utilized the GPT-4o-mini model via the OpenAI API. This model was chosen for its favorable balance of cost, latency, and reasoning quality, enabling scalable evaluation under limited compute budgets. We utilized the following hyperparameter settings for ODR+:
@@
We follow the official BrowseComp evaluation protocol, which requires system responses to conform to a standardized three-part structure (Explanation, Exact Answer, Confidence). Each system output is scored using the released BrowseComp evaluator, which leverages the GPT-4o model (via OpenAI API) to assess both answer correctness and formatting adherence. The evaluator performs semantic comparison between the predicted answer and the ground truth to determine exact match accuracy. Therefore, for each question the evaluator determines whether the response of the DRA is correct or incorrect, and we report the resulting accuracy over the 60 test questions of each system. All web searches and page extractions in ODR and ODR+ were performed using FireCrawl [7] to ensure consistent and structured retrieval.
Main Results
The main results are reported in Table 2, where the results indicate that the ODR system was unable to answer any of the questions in the BrowseComp-Small test set. By contrast, ODR+ answered 10% of test questions (6 of 60) with exact-match correctness. To our knowledge therefore, ODR+ achieves the current state-of-the-art (SOTA) performance on the BrowseComp benchmark, among open-source models. Surprisingly however, we found that ODR+ also outperformed the other two proprietary DRAs that we evaluated on BrowseComp-Small: Claude-DR and Gemini-DR. At the time of our testing, Claude-DR and Gemini-DR both returned 0% exact match accuracy on our test set. In our manual evaluations, we observed that these systems often produced long reports rather than concise answers - an approach that is penalized under BrowseComp’s strict evaluation criteria. By contrast, ODR+ had the advantage that we designed it specifically to solve BrowseComp questions, and therefore it is designed to focus on satisfying the exact constraints of the questions, and generating structured verifiable answers.
For completeness, we also report the results of ChatGPT-DR that were published in [4], which were obtained on the full BrowseComp benchmark. We were unable to evaluate ChatGPT-DR ourselves because it lacked a public API. In [4] the authors reported different performance depending upon the amount of test-time compute that was applied, ranging from approximately 10% with limited test-time compute, up to 51.5% with extensive test-time compute. The precise amount of test-time compute is not provided, making it difficult to judge precisely how this system would compare with ODR+, given the same computational resources. Notably, in contrast to all other systems we consider, the ChatGPT-DR system reported in [4] was developed using the entire BrowseComp benchmark, and without utilizing a disjoint training and testing set, potentially providing it with an advantage.
Table 2: Performance Comparison on BC-Small Test Set
Deep Research Agent LLM Accuracy (%)
ODR GPT-4o-mini 0%
ODR+ (ours) GPT-4o-mini 10%
Claude-DR Sonnet 4 0%
Gemini-DR Gemini 2.5 Pro 0%
ChatGPT-DR GPT-4o $\sim$10–51.5%*
*Results reported from [4] on full BrowseComp Benchmark
Additional Analysis and Ablations for ODR+
In this section we report some additional analysis of ODR+. We first discuss some additional performance metrics for ODR+ that provide insight on its functionality. First we report that ODR+ obtained 20% accuracy on the training set, which is twice as high as we obtained on the testing set. The size of our testing and training set are relatively small, but this could also indicate some degree of overfitting of our system to the training set; this remains an open question. We also found that 25% of test questions (15 of the 60) exceeded the time limit ($T_{max}$), and were terminated before completion, highlighting the potential positive impact of increasing test-time computation. Despite this, 98.3% of all test questions (59/60) produced structurally valid outputs—containing a well-formed Explanation, Exact Answer, and Confidence field—demonstrating strong robustness across system components even in partially completed runs. The average runtime across the 45 successfully completed test questions was 198 seconds (out of a 210-second limit), with longer runtimes typically observed in questions involving more search hops or fallback retries.
@@
  • No Sub-question Decomposition: This variant disables the sub-question generation and decomposition step. The system reverts to a single-query approach similar to the original ODR. We expect a significant drop in performance on multi-hop questions, which typically require breaking down complex prompts into smaller, focused searches.
  • No Iterative Planning: This variant removes adaptive planning and research state management. Sub-questions are processed sequentially, without retry logic or feedback based on earlier findings. This limits the system’s ability to adjust its strategy dynamically, likely reducing the effectiveness of information gathering.
  • No Structured Synthesis: This variant disables structured output formatting and validation. Instead, it uses free-text generation similar to the original ODR. We anticipate lower confidence estimates, formatting inconsistencies, and a higher likelihood of inaccurate or incomplete final answers.
@@ 7  Conclusions
We introduced ODR+, an enhanced open-source Deep Research Agent (DRA) designed to perform complex multi-hop web-based question answering. Building on the original ODR system, ODR+ incorporates sub-question decomposition, iterative planning, and structured synthesis, enabling more effective and interpretable autonomous research across open-domain queries.
We benchmarked ODR+ on the BrowseComp-Small dataset and demonstrated that it significantly outperforms the original ODR baseline, achieving 10% exact-match accuracy on the test set while maintaining high structural validity. Our ablation studies confirmed the critical role of each added component in improving performance.
To support continued progress in the development and evaluation of DRAs, we release our implementation and tools publicly. We hope ODR+ serves as a foundation for future research in open, analyzable, and extensible Deep Research Agents.

### v2

@@ 5.2 Iterative Sub-Solution Search
This module focuses on addressing each of the sub-questions identified in module (1) and is shown in lines 10–30 of Algorithm 1. The iterative process continues until all sub-questions are addressed, or some other stopping criteria is met (e.g., permissible run-time, denoted $T_{max}$, is exceeded; or a maximum number of sub-questions, denoted $D_{max}$ is exceeded).
@@
At the beginning of each iteration, the system selects an unresolved sub-question from S.subquestions (Line 12) and uses it directly as a web search query. We observed that web browsers returned a different page ranking each time the same query was submitted and therefore we submitted the same query $N_{\text{query}}$ times using webSearch (Line 13). The top-ranked URLs are gathered from each of the $N_{\text{query}}$ searches. The $k$ most frequently occurring URLs are then chosen for additional processing using selectMostFrequent (Line 14) after the frequency of each URL across all attempts is totaled.
@@
Then, by calling createExtractionPrompt with the prompt template $P_{3}$, $userQuery$, and the constraints that were previously extracted in module (1), an extraction prompt extractionPrompt is created (Line 15). It is intended to give the LLM instructions to extract only the parts of the page content that are relevant to addressing the sub-question. The extractionPrompt is passed to an LLM, along with the full text of each selected URL. The LLM is invoked once per URL and typically returns one, or a few, short spans of relevant text. These outputs are stored as structured findings—each consisting of the extracted text and its source URL—and are appended to S.findings, the list of accumulated findings maintained in the internal research state (Line 17).
@@
After collecting new findings, ODR+ invokes an LLM using analyzeEvidence with prompt $P_{4}$, the current sub-question, and the full set of accumulated findings (Line 18). The prompt directs the model to generate a structured response in JSON format, which includes fields like a confidence score, a list of satisfied constraints, a proposed answer to the sub-question (if one can be found), and any recommended follow-up sub-questions. A valid response is added to S.subAnswers (Lines 19–20) following the parsing of the JSON output. If follow-up subquestions are suggested, they are added to S.subquestions (Lines 21–22). The model’s analysis, particularly the confidence score and recommendation on whether to proceed, is also used by the control flow logic to decide whether to proceed to the next iteration or terminate the loop early (Lines 23–24).
@@
5.3 Response Synthesis
The third and final module of the ODR+ system, which is implemented in lines 32–33 of Algorithm 1, is responsible for synthesizing the final structured answer. The system uses the engineered prompt $P_{5}$ to invoke an LLM on line 32. It also includes the original user question (userQuery), the extracted constraints (constraints), and the accumulated evidence (S.findings). The prompt specifically prohibits reliance on prior knowledge and directs the model to produce a final response based only on this structured content. The model is asked to produce a response in the standardized BrowseComp format:
@@ 5.3 Response Synthesis
The prompt also instructs the model to compute a confidence score based on the number of key constraints satisfied by the proposed answer, relative to the total number of extracted constraints. On line 33, the model’s output is stored in structuredResponse. The system then validates this response to ensure that all required fields are present and correctly formatted. If any field (such as the explanation, exact answer, or confidence score) is missing or malformed, fallback values are inserted. For example, the system may assign "Unknown" as the answer and a default confidence score of 10%. This validation step ensures that every final output is complete, properly structured, and ready for automated evaluation.
@@ 6 Numerical Experiments
We conduct experiments on our BrowseComp-Small benchmark (see Sec. 3), which comprises two disjoint sets of sixty questions: a training set and a testing set. We evaluate several competing DRA systems on the sixty test questions: ODR, ODR+, Claude-DeepResearch (Anthropic), and Gemini-DeepResearch (Google 2025).
6.1 ODR+ Development and Hyperparameter Settings
All development of ODR+ was done using the sixty training questions in our BrowseComp-Small benchmark. This was done to minimize the potential of overfitting the design of ODR+ to the testing questions. Many steps of ODR+ (and ODR) utilize an LLM, and we utilized the GPT-4o-mini model via the OpenAI API. This model was selected because it allows for scalable evaluation under constrained compute budgets and offers a good trade-off between cost, latency, and reasoning quality. For ODR+, we used the following hyperparameter settings:
@@ 6.2 Evaluation Methodology
We follow the official BrowseComp evaluation protocol, which requires system responses to conform to a standardized three-part structure (Explanation, Exact Answer, Confidence). Each system output is scored using the released BrowseComp evaluator, which leverages the GPT-4o model (via OpenAI API) to assess both answer correctness and formatting adherence. The evaluator performs semantic comparison between the predicted answer and the ground truth to determine exact match accuracy. Therefore, for each question the evaluator determines whether the response of the DRA is correct or incorrect, and we report the resulting accuracy over the 60 test questions of each system. All web searches and page extractions in ODR and ODR+ were performed using FireCrawl to ensure consistent and structured retrieval.
6.3 Main Results
The main results are reported in Table 2. ODR was unable to answer any questions in the BrowseComp-Small test set, whereas ODR+ answered 10% (6 of 60) with exact-match correctness. In BrowseComp, “exact match” is determined by the official evaluator, which requires structured responses (Explanation, Exact Answer, Confidence) and uses GPT-4o to check semantic equivalence with the ground truth. Because BrowseComp answers are short (e.g., names, numbers, or short phrases), this evaluation is highly reliable. To our knowledge, ODR+ achieves the current state-of-the-art (SOTA) performance on the BrowseComp benchmark among open-source models.
Surprisingly, ODR+ also outperformed the two proprietary DRAs we tested: Claude-DR and Gemini-DR, both of which achieved 0% accuracy on the 60-question test set. Because these systems do not expose structured outputs, we manually reviewed their answers against the ground truth. In nearly all cases, their outputs were long, report-style responses rather than the short exact answers required by the benchmark. We inspected these generated report, and confirmed that they did not contain the correct answers, so their accuracy remained 0%. We note that ODR+ was developed using a separate 60-question training split, whereas ODR, Claude-DR, and Gemini-DR were evaluated zero-shot on the test set, introducing a potentially significant disadvantage for them. Unfortunately at the time of our experimentation, these proprietary systems could not be tuned or adjusted for a custom benchmark such as BrowseComp. Our experiments represent our best attempt to evaluate them fairly and transparently, however, our methods still imposed the aforementioned disadvantages.
For completeness, we also report the results of ChatGPT-DR that were reported in [10], which were obtained on the full BrowseComp benchmark, and which varied depending on test-time compute, from $\sim$10% with limited compute to 51.5% with extensive compute. The paper shows performance scaling with browsing effort and sampling, but does not specify the exact compute allocations for these settings, making direct comparison difficult. Unlike our setup, ChatGPT-DR was potentially developed using the entire BrowseComp benchmark rather than a disjoint train/test split, which may provide an advantage.
In addition to accuracy, we also measured average wall-clock runtime. ODR+ required $\sim$198 seconds per question, close to its fixed 210 s limit. This time limit kept bounded compute, and ODR+ typically used the full available budget. By contrast, ODR failed to complete runs, while Claude-DR averaged  11 minutes and Gemini-DR  4 minutes per question under default APIs. Thus, ODR+’s performance cannot be attributed to greater compute availability, since proprietary systems actually consumed more time on average.
Table 2: Performance and Runtime Comparison on BC-Small Test Set
Deep Research Agent LLM Accuracy (%) Avg. Runtime / Q
ODR GPT-4o-mini 0% N/A
ODR+ (ours) GPT-4o-mini 10% 198s
Claude-DR Sonnet 4 0% 11 min
Gemini-DR Gemini 2.5 Pro 0% 4 min
ChatGPT-DR GPT-4o $\sim$10–51.5%* N/A
*Results reported from [10] on the full BrowseComp benchmark.
6.4 Ablation Studies
@@
  • No Sub-question Decomposition: The sub-question generation and decomposition step is disabled in this variant. The system returns to the original ODR’s single-query methodology. On multi-hop questions, which usually call for breaking down complex prompts into more manageable, targeted searches, we anticipate a notable decline in performance.
  • No Iterative Planning: Adaptive planning and research state management are eliminated in this variant. Sub-questions are handled one after the other without the use of retry logic or feedback based on past results. This restricts the system’s ability to dynamically modify its approach, which probably lowers the efficiency of information gathering.
  • No Structured Synthesis: This variant eliminates structured output formatting and validation. It uses free-text generation like the original ODR instead. We expect lower estimates of confidence, formatting problems, and a higher chance of getting final answers that are wrong or incomplete.
@@ 7 Conclusions
We introduced ODR+, an enhanced open-source Deep Research Agent (DRA) designed to perform complex multi-hop web-based question answering. Building on the original ODR system - the only open-source DRA we could identify at the outset of this research - ODR+ incorporates several improvements: sub-question decomposition, iterative planning, and structured synthesis. We benchmarked ODR+ on the BrowseComp-Small dataset, a subset of the BrowseComp benchmark that we curated for more scalable DRA benchmarking, and demonstrated that it significantly outperforms the original ODR baseline, achieving 10% exact-match accuracy on the test set while producing answers in the required format (Explanation, Exact Answer, Confidence). We also present evidence that ODR+ is competitive with proprietary systems, although fair comparisons are difficult. Our ablation studies confirmed the critical role of our three proposed improvements over ODR. To support continued progress in the development and evaluation of DRAs, we release our implementation and tools publicly. We hope ODR+ serves as a foundation for future research in open, analyzable, and extensible Deep Research Agents.